OpenStudy (anonymous):

Given the equation 2x + 4 = 4x - 2, select the reasoning that correctly solves for x.

OpenStudy (anonymous):

@jacalneaila help last one plz

OpenStudy (anonymous):

@uri

OpenStudy (uri):

Do you want the answer or do you want to learn?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

2x+4-4x-2=0 I think.

OpenStudy (uri):

I will help you,You need to find the answer yourself @350zlove

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I gave you the beginning, you finish it O.K. ?

OpenStudy (anonymous):

subtract 4x from both sides. then subtract 4 from both sides then what do you have ?

OpenStudy (anonymous):

Add 2, subtract 2x, then divide by 2. Add 2, subtract 4x, then divide by -2. Subtract 4, subtract 2x, then divide by -2. Subtract 4, subtract 4x, then divide by 2. Add 2, subtract 2x, then divide by 2 Add 2, subtract 4x, then divide by -2.

OpenStudy (uri):

2x + 4 = 4x - 2, Subract 4 both sides 2x + 4 = 4x - 2 -4 -4 -------------- What is left?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

@uri 2

OpenStudy (anonymous):

x=3 thats the answer ... i gave u the answer now try to solve it

OpenStudy (anonymous):

thats correct not 2

OpenStudy (anonymous):

combine the like terms

OpenStudy (anonymous):

i will show the solution for the last time ok @350zlove

OpenStudy (anonymous):

2x+4=4x-2 2x-4x=-2-4 -2x= -6

OpenStudy (anonymous):